If v(1) = < a,-b > and v(2) = < -2a, b > does v(1)+v(2) = 0?
Answer by Fombitz(32388)

V1=(a,-b)
V2=(-2a,b)
.
.
V1+V2=(a-2a,-b+b)
V1+V2=(-a,0)
No, it does not equal (0,0).